# Pushing the frontiers of subexponential FPT time for Feedback Vertex Set

### v1

@@ 3.8. Final step when $G-M$ admits a $t$-uniform partition with a small number of large-degree parts
As $A(G,k,M,\mathcal{H})$ reaches Line 23, it implies that Partitionning-Algorithm  output a $t$-uniform partition of $G-M$ (where $t=2d_{r}$ and $|\mathcal{T}^{+}|\leq p_{3}(r,s,t)|M|$). Let $\tilde{M}=M\cup Z_{1}(\mathcal{T})\cup T_{2}(\mathcal{T})$. By 3.31, Rule (KR1) and Rule (KR1) do not apply on $(G,k,\tilde{M},\mathcal{H})$. As we did not apply Rule (KR4) (Line 16) or Rule (KR5) (Line 19), and as by 3.31, $d_{\tilde{M}}(T)\leq 2t+2$ for any connected component $T$ of $G-\tilde{M}$, we can apply 3.32 to obtain $|G-\tilde{M}|\leq p_{4}(r,t)|\tilde{M}|$. As $|G-M|=|G-\tilde{M}|+|Z_{1}(\mathcal{T})|+|Z_{2}(\mathcal{T})|$, we get $|G-M|=\mathcal{O}\left(p_{4}(r,t)(|M|+|Z_{1}(\mathcal{T})|+|Z_{2}(\mathcal{T})|)\right)$. By 3.31 we have $|Z_{1}(\mathcal{T})|+|Z_{2}(\mathcal{T})|\leq 2|\mathcal{T}^{+}|$ and $|\mathcal{T}^{+}|\leq p_{3}(r,s,t)|M|$, we get $|G-M|=\mathcal{O}(p_{3}(r,t)p_{4}(r,t)|M|)$. ∎
Let us now bound the size of $M$ in any call to $A(G,k,M,\mathcal{H})$ by providing invariants on the input of Algorithm 1. Informally, $|M|$ can be bounded as follows. Recall that we perform our first call with $A(G_{i},k_{i},M_{i},\emptyset)$ where $(G_{i},k_{i},M_{i})\in\mathcal{I}$ and $M_{i}\subseteq M_{0}$, with $|M_{0}|\leq 2k_{0}$. Then, if we follow any path in the tree of calls whose root is $A(G_{i},k_{i},M_{i},\emptyset)$, we may add some vertices to $M$ because of (KR3), let denote $A_{1}$ those vertices, or because of (BR2), we denote $A_{2}$ those vertices. On a given branch the rule (BR2) add at most $k_{0}$ trees of size at most $p_{2}(r,t)$ so $|A_{2}|\leq k_{0}p_{2}(r,t)$. Remains to bound the vertices of $A_{1}$. Observe that when a vertex $v$ is added to the set $M$ because of (KR3) (and so to $A_{2}$) we have $d_{M}(v)\geq 2d_{r}$, and so the number of edges in the graph $G[M]$ increases by at least $2d_{r}$. We may think that as we have at most $d_{r}|M|$ edges in $G[M]$, we obtain $2d_{r}|A_{2}|\leq d_{r}|M|\leq d_{r}(|M_{0}|+|A_{1}|+|A_{2}|)$ which gives a bound on the size of $A_{1}$. However this reasoning is flawed as we needed to consider all the vertices added to $M$, even the one deleted because of (KR5). But then in the supergraph of $G^{\prime}$ containing those vertices they may be some vertices obtained after contraction of edges by (KR2), which were not of degree $2$ in their current supergraph considering deleted vertices. So this supergraph may not be in $\mathcal{G}$, and in fact may even contains a $K_{r,r}$ as subgraph. So a more technical analysis is needed to bound the size of $M$.
@@
For any input $(G,k,M,\mathcal{H})$ of Algorithm 1, we have $|M|=\mathcal{O}(k_{0}p_{5}(r,t))$ with $p_{5}(r,t)=4+2p_{2}(r,t)$.
@@
For proving the result, we need to define some invariant to our problem. For stating this invariant we need to consider additional annotations to our problem. We now consider instances
$(G,k,M,\mathcal{H},A_{0},A_{1},A_{2},X,C,G^{\prime})$. Informally $A_{0}$ will denote the vertices of $M\cap M_{0}$, $A_{1}$ the vertices of $M$ added by (KR3), $A_{2}$ for the ones added to $M$ by rule (BR2), $C$ the set of vertices obtained by contracting at least one edge by (KR2), $X$ the vertices removed by application of (KR5) (so $X$ is a set of vertices outside of $V(G)$), and $G^{\prime}$ is a supergraph of $G$ with $V(G^{\prime})=(V(G)\cup X)$ that will be useful to “remember” the links between the vertices of $V(G)$ and $X$. More formally, initially for $(G,M,k,\emptyset)\in\mathcal{I}$ the instance is such that $A_{1}=M$, $A_{2}=A_{3}=X=C=\emptyset$ and $G^{\prime}=G$. The generated instance update the additional annotations in the following manner:
  • When removing vertices from $V(G-M)$, so for (KR1) and (KR4), remove the vertices from $G^{\prime}$ as well and from the sets containing them.
@@
  • If this is (KR3) which is applied, the vertex $v$ added to $M$ is added to $A_{1}$ as well.
  • For (KR5), the vertex $v\in M$ removed from $G$ is now added to $X$, and removed from the sets $A_{i}$ in which it is. $G^{\prime}$ is not modified.
  • For (BR2), in the $t$ branches where we add vertices to $M$, we add them to $A_{2}$ as well now. For the $t$ branches where $t-1$ vertices are removed from $M$, again there are now added to $X$, and removed from the sets $A_{i}$ which contain them.
@@
The first claim is trivial to verify. For the second, observe that the only modification to this graph is the deletion of vertices, so the result follows by hereditary of $\mathcal{G}$. For each call the degree of a vertex does not increases in $G$ at any time (but it can in $G^{\prime}$), so it suffice to prove the result when the vertex was just added. If a vertex $v\in V(G)$ is added to $C$ or one of its neighbor it implies that $d_{G}(v)\leq 2$. Finally, for the last claim, this property is true when a vertex is added to $A_{2}$ and there is nor contraction of edges or deletion of vertices in $G^{\prime}[M\cup X]$, so the inequality is kept. The fourth result can be proved by observing that an element of the intersection would have first been in $C$ and then $A_{1}$. Adding an element in $A_{1}$ requires that it has $22d_{r}>$ 2 neighbors in $G$, but this would contradict the previous bullet. Finally the last inequality is obtained by observing that the considered set stays the same once $v$ is added to $A_{1}$. And when $v$ is added to $A_{1}$ we have $d_{M}(v)\geq 2d_{r}>2$, so $N_{G^{\prime}}(v)\cap C=\emptyset$ and so $N_{G^{\prime}}(v)\subseteq(M\cup X)\setminus C$
Now following the same reasoning than in the sketch but this time using the graph $G^{\prime}[(V(G)\cup X)\setminus C]$, we obtain $|A_{0}|\leq 2k_{0}$, $|A_{2}|\leq k_{0}p_{2}(r,t)$ and $2d_{r}|A_{1}|\leq d_{r}(|A_{0}|+|A_{1}|+|A_{2}|)$ so $|M|\leq 4k_{0}+2k_{0}p_{2}(r,t)$. ∎
Using 3.33 and 3.34, the following corollary is now immediate.
@@
Let $(G,k,M,\mathcal{H})$ be an input of $(r,\mathcal{G})$-Ann-FVS. If $A(G,k,M,\mathcal{H})$ reaches Line 23, then $|V(G)|=\mathcal{O}(k_{0}p_{6}(r,t))$ where $p_{6}(r,t)=p_{3}(r,t)p_{4}(r,t)p_{5}(r,t)$.
@@ 4. Complexity analysis
4.1. Complexity of the DP in the base case of Algorithm 1
Let us first describe how Algorithm 1 solves an instance $(G,k,M,\mathcal{H})$ using dynamic programming when no rule applies (DP in the Line 23). Recall that by definition of $(r,\mathcal{G})$-Ann-FVS, $G$ is a $K_{r,r}$-free graph from a nice graph class $\mathcal{G}$, and $\mathcal{H}$ a family of disjoint subsets of $M$, with each $H\in\mathcal{H}$ inducing a connected graph in $G$.
Theorem 4.1 ([CFK+15]).
@@ 4.1. Complexity of the DP in the base case of Algorithm 1
For an instance $(G,M,k,\mathcal{H})$ with $G$ a graph on $n$ vertices, computing a tree decomposition with bags of size $\operatorname{\textsf{tw}}(G)$ up to a constant factor can be done in time $2^{\mathcal{O}(\operatorname{\textsf{tw}}(G))}n$ [Kor22]. Given such a tree decomposition observe that a bag of size $b$ can intersect at most $b$ distinct $H\in\mathcal{H}$. Moreover for $H\in\mathcal{H}$, having that $H$ is connected ensures that the set of bags of the tree decomposition of $G$ containing at least one vertex of $H$ is a connected subset. Using those two observations it is easy to adapt the standard dynamic programming algorithm solving FVS for solving the more general $(r,\mathcal{G})$-Ann-FVS without worsening the running time, and obtaining the claimed complexity. ∎
@@
Let $(G,k,M,\mathcal{H})$ be an input of $(r,\mathcal{G})$-Ann-FVS. If $A(G,k,M,\mathcal{H})$ reaches Line 23 and calls $DP(G,k,M,\mathcal{H})$, then the worst case running time of $DP(G,k,M,\mathcal{H})$ is $T_{DP}(k_{0},r,t)=2^{\mathcal{O}\left(k_{0}^{\delta}p_{7}(r,t)\right)}$ with $p_{7}(r,t)=f_{r}\log(f_{r}k_{0}p_{6}(r,t))(p_{6}(r,t))^{\delta}$.
@@
Denoting $n=|V(G)|$, according to 3.36, we have $n=\mathcal{O}(k_{0}p_{6}(r,t))$, and as $\mathcal{G}$ is a nice class (see 1.4), we have $\operatorname{\textsf{tw}}(G)\leq f_{r}n^{\delta}$. Thus, Theorem 4.1 implies that $DP(G,k,M,\mathcal{H})$ runs in time $\left(f_{r}n\right)^{\mathcal{O}\left(f_{r}n^{\delta}\right)}n^{\mathcal{O}(1)}$. ∎
@@ 4.2. Complexity of Algorithm 1
Informally, the complexity of Algorithm 1 is dominated by the only branching rule ((BR2)). The appropriate parameter $\alpha$ associated to an instance $(G,k,M,\mathcal{H})$ is $\alpha=k+(k-|\mathcal{H}|)$, as informally $\mathcal{H}$ is a packing of hyperedges that we have to hit, and thus $|\mathcal{H}|$ is a lower bound of the cost of any solution. Thus, if $p(n)$ is the (polynomial) complexity of all operations performed in one call of Algorithm 1, and $f(n,\alpha)$ is the worst complexity of Algorithm 1 when $|V(G)|=n$ and $\alpha=k+(k-|\mathcal{H}|)$, (BR2) leads to $f(n,\alpha)\leq p(n)+(2t)f(n,\alpha-(t-1))$, leading to $f(n,\alpha)\leq p(n)(2t)^{\frac{2k-h}{t-1}}$.
@@
Given an input $(G,k,M,\mathcal{H})\in\mathcal{I}$ with $G$ a $n$-vertex graph, $A(G,k,M,\mathcal{H})$ runs in time
@@
Let $f(n,k,m,h)$ be the worst case complexity of $A(G,k,M,\mathcal{H})$ when $|V(G)|=n$, $|M|=m$, and $\mathcal{H}=h$. We denote $C_{DP}=T_{DP}\left(k_{0},r,t\right)$. Let $p(n)$ be the worst case complexity of all operations performed in one call of Algorithm 1 (which is bounded by the sum of the complexity of testing and applying all the kernelization and branching rules, and applying Subsection 3.7). Notice that, as all rules and Subsection 3.7 are polynomial, $p$ is polynomial. Let us show by induction (on $2(n+k)-m-h$) that
@@
Without loss of generality, we assume that $p$ is non-decreasing. Let us bound $f$ according to which recursive call we perform. In particular, if we apply (KR5), we delete one vertex $v$ from $M$, and as we remove from the packing $\mathcal{H}$ the (at most one) hyperedge containing $v$, we decrease $h$ by $x\in\{0,1\}$. If we apply (BR2), then we make $t$ recursive calls on instances $((G-X_{\overline{v}}),k-(t-1),M\setminus X_{\overline{v}},\mathcal{H}-X_{\overline{v}})$ (re-using notations of (BR2)), and thus in these calls we remove $x\in[t-1]$ hyperedges from $\mathcal{H}$, and in the worst case also $t$ recursive calls on instances $\left(G,k,M\cup R_{i},\mathcal{H}\cup\mathcal{T}_{R_{i}}\right)$, where $R_{i}$ is the union of $t-1$ parts of a $\widetilde{K_{t,t}}$, each part having size at most $p_{2}(r,t)$ by definition of a $\widetilde{K_{t,t}}$. This leads to the following bound:
[…]
It remains to check by induction that our bound hold in any of these cases. The first two and last cases are straightforward. For the third case, we get
@@
This ends the induction. The wanted result is then easily obtained from the proven bound. ∎
@@
Let denote $n_{0}=|V(G_{0})|$. According to 3.4, the complexity to generate the set $\mathcal{I}$ is in $2^{\mathcal{O}\left(r\log k_{0}\right)}n_{0}^{\mathcal{O}(1)}(2r)^{\frac{k_{0}}{r-1}}$ and $|\mathcal{I}|=\mathcal{O}\left((2r)^{\frac{k_{0}}{r-1}}\right)$. Deciding one instance $(G,k,M,\mathcal{H})\in\mathcal{I}$ with $G$ a $n$-vertex graph can be done in time $T_{DP}(k_{0},r,t)(2t)^{\frac{2k_{0}}{t-1}}n_{0}^{\mathcal{O}(1)}$ where $T_{DP}(k_{0},r,t)=2^{\mathcal{O}\left(k_{0}^{\delta}p_{7}(r,t)\right)}$ according to 4.3, 4.2 and the observation that $k\leq k_{0}$ and $n\leq n_{0}$.
@@
Now lets recall that $t$ is defined in function of $r$ as $t=2d_{r}=r^{\mathcal{O}(1)}$ and so $p_{7}(r,t)=r^{\mathcal{O}(1)}$. More precisely if we have $f_{1}(r)=\operatorname{\tilde{\mathcal{O}}}\left(r^{c_{f_{1}}}\right),~{}f_{2}(r,p,m)=\operatorname{\tilde{\mathcal{O}}}(r^{c_{f_{2}}}(p+m)^{c_{f_{2}}^{\prime}})$, $f_{r}=\operatorname{\tilde{\mathcal{O}}}(r^{c_{f}})$ and $d_{r}=\operatorname{\tilde{\mathcal{O}}}(r^{c_{d}})$ (remember that without loss of generality we supposed $c_{d}\geq 1$), then we have $p_{7}(r,t)=\operatorname{\tilde{\mathcal{O}}}(r^{c_{7}})$ with
@@
So by taking $r=k_{0}^{\varepsilon}$ with $\varepsilon=\frac{1-\delta}{c_{7}+1}$ we obtain a running time $2^{\operatorname{\tilde{\mathcal{O}}}\left(k_{0}^{1-\varepsilon}\right)}n_{0}^{\mathcal{O}(1)}$. To get rid of the hidden logarithmic factors in the exponent and obtain the wanted result it then suffices to replace $\varepsilon$ with $0<\varepsilon^{\prime}<\varepsilon$. ∎

### v2

@@ 3.8. Final step when GMG-M admits a $t$-uniform partition with a small number of large-degree parts
[…]
Let us now bound the size of $M$ in any call to Algorithm 1 by providing invariants on its inputs. Informally, $|M|$ can be bounded as follows. Recall that we perform our first call to Algorithm 1 on input $(G_{i},k_{i},M_{i},\emptyset)$ where $(G_{i},k_{i},M_{i})\in\mathcal{I}$ and $M_{i}\subseteq M_{0}$, with $|M_{0}|\leq 2k_{0}$. Then, if we follow any path in the tree of calls whose root is the call on input $(G_{i},k_{i},M_{i},\emptyset)$, we may add some vertices to $M$ because of (KR3) (let denote $A_{1}$ those vertices) or because of (BR2) (we denote $A_{2}$ those vertices). On a given branch the rule (BR2) add at most $k_{0}$ trees of size at most $p_{2}(r,t)$ so $|A_{2}|\leq k_{0}p_{2}(r,t)$. It remains to bound the vertices of $A_{1}$. Observe that when a vertex $v$ is added to the set $M$ because of (KR3) (and so to $A_{2}$) we have $d_{M}(v)\geq 2d(r)$, and so the number of edges in the graph $G[M]$ increases by at least $2d(r)$. We may think that as we have at most $d(r)|M|$ edges in $G[M]$, we obtain $2d(r)|A_{1}|\leq d(r)|M|\leq d(r)(|M_{0}|+|A_{1}|+|A_{2}|)$ which gives a bound on the size of $A_{1}$. However this reasoning is flawed as we actually need to consider all the vertices added to $M$, even those deleted because of (KR5). But then, if we consider deleted vertices then some vertices which were handled by (KR2) may not have degree 2 anymore, and the graph we need to consider possibly does not live in $\mathcal{G}$, and in fact may even contains a $K_{r,r}$ as a subgraph. Therefore a more technical analysis is needed to bound the size of $M$.
@@
In a run of Algorithm 1 on some input $(G_{i},k_{i},M_{i},\emptyset)$, where $(G_{i},k_{i},M_{i})\in\mathcal{I}$, there may be several calls to Algorithm 1, from the initial one to possible recursive calls. Any such call is made on an input of the form $(G,k,M,\mathcal{H})$ with $|M|=\mathcal{O}(k_{0}p_{2}(r,t))$.
@@
For proving the result, we need to define some invariant to our problem. For stating this invariant we need to consider additional annotations to our problem. We now consider instances of the form
$(G,k,M,\mathcal{H},A_{0},A_{1},A_{2},X,C,G^{\prime}).$
Informally $A_{0}$ will denote the vertices in $M$ from the beginning, $A_{0}=M\cap M_{0}$, $A_{1}$ the vertices of $M$ added by (KR3), $A_{2}$ for the ones added to $M$ by the rule (BR2), $C$ the set of vertices obtained by contracting at least one edge by (KR2), $X$ the vertices removed by application of (KR5) (so $X$ is a set of vertices outside of $V(G)$), and $G^{\prime}$ is a supergraph of $G$ with $V(G^{\prime})=(V(G)\cup X)$ that will be useful to “remember” the links between the vertices of $V(G)$ and $X$. More formally, initially for $(G,M,k,\emptyset)\in\mathcal{I}$ the instance is such that $A_{0}=M$, $A_{1}=A_{2}=X=C=\emptyset$ and $G^{\prime}=G$. The rules update the additional annotations in the following manner:
  • When removing vertices from V(GM)V(G-M) in (KR1) and (KR4), remove the vertices from $G^{\prime}$ as well and from the sets containing them.
@@
  • If we apply (KR3), the vertex $v$ added to $M$ is added to $A_{1}$ as well.
  • If we apply (KR5), the vertex $v\in M$ removed from $G$ is now added to $X$, and removed from the set $A_{0},A_{1}$, or $A_{2}$ it belongs to. The graph $G^{\prime}$ is not modified.
  • If we apply (BR2), in the $t$ branches where we add vertices to $M$, we add them to $A_{2}$ as well now. For the $t$ branches where t1t-1 vertices are removed from $M$, again they are now added to $X$, and removed the sets $A_{0},A_{1}$, and $A_{2}$.
@@
[…]
Now following the same reasoning than in the sketch but this time using the graph $G^{\prime}[(V(G)\cup X)\setminus C]$, we obtain $|A_{0}|\leq 2k_{0}$, $|A_{2}|\leq k_{0}p_{2}(r,t)$ and $2d(r)|A_{1}|\leq d(r)(|A_{0}|+|A_{1}|+|A_{2}|)$ so $|M|\leq 4k_{0}+2k_{0}p_{2}(r,t)$. ∎
Using 3.31 and 3.32, the following corollary is now immediate.
@@
Let $(G,k,M,\mathcal{H})$ be an instance of $(r,\mathcal{G})$-Ann-FVS. If on this input Algorithm 1 reaches line 23, then $|V(G)|=\mathcal{O}(k_{0}p_{6}(r,t))$ where $p_{6}(r,t)=p_{2}(r,t)p_{3}(r,t)p_{4}(r,t)$.
@@ 4. Complexity analysis
4.1. Complexity of the dynamic programming in the base case of Algorithm 1
Let us first describe how Algorithm 1 solves an instance $(G,k,M,\mathcal{H})$ using dynamic programming when no rule applies (line 23). Recall that by definition of $(r,\mathcal{G})$-Ann-FVS, $G$ is a $K_{r,r}$-free graph from a nice graph class $\mathcal{G}$, and $\mathcal{H}$ a family of disjoint subsets of $M$, with each $H\in\mathcal{H}$ inducing a connected graph in $G$.
Theorem 4.1.
@@ 4.1. Complexity of the dynamic programming in the base case of Algorithm 1
For an instance $(G,M,k,\mathcal{H})$ with $G$ a graph on $n$ vertices, computing a tree decomposition with bags of size $\mathcal{O}(\operatorname{\textsf{tw}}(G))$ can be done in time $2^{\mathcal{O}(\operatorname{\textsf{tw}}(G))}n$ [korhonen2022single]. Given such a tree decomposition, observe that a bag of size $b$ can intersect at most $b$ distinct $H\in\mathcal{H}$. Moreover, for $H\in\mathcal{H}$, the fact that $H$ is connected ensures that the set of bags of the tree decomposition of $G$ containing at least one vertex of $H$ is a connected subset. Using those two observations it is straightforward to adapt the standard dynamic programming algorithm solving FVS (see [Cygan2015Book]) for solving the more general $(r,\mathcal{G})$-Ann-FVS without worsening the running time, and obtaining the claimed complexity. ∎
Let us call DP the algorithm of Theorem 4.1.
@@
Let $(G,k,M,\mathcal{H})$ be an instance of $(r,\mathcal{G})$-Ann-FVS. If on this instance Algorithm 1 reaches line 23 and calls $DP(G,k,M,\mathcal{H})$, then the worst case running time of $DP(G,k,M,\mathcal{H})$ is $T_{DP}(k_{0},r,t)=2^{\mathcal{O}\left(k_{0}^{\delta}p_{7}(r,t)\right)}$ with $p_{7}(r,t)=\tau(r)\log(\tau(r)k_{0}p_{6}(r,t))(p_{6}(r,t))^{\delta}$.
@@
Denoting $n=|V(G)|$, according to 3.34, we have $n=\mathcal{O}(k_{0}p_{6}(r,t))$, and as $\mathcal{G}$ is a nice class (see 1.4), we have $\operatorname{\textsf{tw}}(G)\leq\tau(r)n^{\delta}$. Thus, Theorem 4.1 implies that $DP(G,k,M,\mathcal{H})$ runs in time $\left(\tau(r)n\right)^{\mathcal{O}\left(\tau(r)n^{\delta}\right)}n^{\mathcal{O}(1)}$. ∎
@@ 4.2. Complexity of Algorithm 1
Informally, the complexity of Algorithm 1 is dominated by the only branching rule (BR2). The appropriate parameter $\alpha$ associated to an instance $(G,k,M,\mathcal{H})$ is $\alpha=k+(k-|\mathcal{H}|)$, as informally $\mathcal{H}$ is a packing of hyperedges that we have to hit, and thus $|\mathcal{H}|$ is a lower bound of the cost of any solution. Thus, if $p(n)$ is the (polynomial) complexity of all operations performed in one call of Algorithm 1, and $f(n,\alpha)$ is the worst complexity of Algorithm 1 when $|V(G)|=n$ and $\alpha=k+(k-|\mathcal{H}|)$, (BR2) gives $f(n,\alpha)\leq p(n)+(2t)f(n,\alpha-(t-1))$, leading to $f(n,\alpha)\leq p(n)(2t)^{\frac{2k-h}{t-1}}$.
@@
On an input $(G,k,M,\mathcal{H})\in\mathcal{I}$ with $G$ a $n$-vertex graph, Algorithm 1 runs in time
@@
Let $f(n,k,m,h)$ be the worst case complexity of Algorithm 1 on an instance $(G,k,M,\mathcal{H})$ when $|V(G)|=n$, $|M|=m$, and $\mathcal{H}=h$. We set $C_{DP}=T_{DP}\left(k_{0},r,t\right)$. Let $p(n)$ be the worst case complexity of all operations performed in one call of Algorithm 1 (which is bounded by the sum of the complexities of testing and applying all the kernelization and branching rules, and applying Partitionning-Algorithm  ). Notice that, as all rules and Subsection 3.7 are polynomial, $p$ is polynomial. Let us show by induction on 2(n+k)mh2(n+k)-m-h that
@@
Without loss of generality, we assume that $p$ is non-decreasing. Let us bound $f$ according to which recursive call we perform. In particular, if we apply (KR5), we delete one vertex $v$ from $M$, and as we remove from the packing $\mathcal{H}$ the (at most one) hyperedge containing $v$, we decrease $h$ by $x\in\{0,1\}$. If we apply (BR2), then we make $t$ recursive calls on instances $((G-X_{\overline{v}}),k-(t-1),M\setminus X_{\overline{v}},\mathcal{H}-X_{\overline{v}})$ (re-using notations of (BR2)), and thus in these calls we remove $x\in\left\{0,\dots,t-1\right\}$ hyperedges from $\mathcal{H}$, and in the worst case also $t$ recursive calls on instances $\left(G,k,M\cup R_{i},\mathcal{H}\cup\mathcal{T}_{R_{i}}\right)$, where $R_{i}$ is the union of t1t-1 parts of a $\widetilde{K_{t,t}}$, each part having size at most $p_{2}(r,t)$ by definition of a $\widetilde{K_{t,t}}$. Thus $f(n,k,m,h)$ is at most:
  1. (1)
    p(n)+f(nx,k,m,h)p(n)+f(n-x,k,m,h), for $x\geq 1$, when (KR1), (KR2), or (KR4) applies;
  2. (2)
    $p(n)+f(n,k,m+1,h)$ when (KR3) applies;
  3. (3)
    p(n)+f(n1,k1,m1,hx)p(n)+f(n-1,k-1,m-1,h-x) for $x\in\{0,1\}$ when (KR5) applies;
  4. (4)
    $p(n)+tf(n-(t-1),k-(t-1),m-(t-1),h-x)+tf(n,k,m+(t-1)p_{2}(r,t),h+(t-1))$ for $0\leq x\leq t-1$ when (BR2) applies; and
  5. (5)
    $p(n)+C_{DP}$ if we do not apply any rule and call the dynamic programming algorithm DP.
It remains to check by induction that our bound holds in any of these cases. The first two and last cases are straightforward. For the third case, we get
@@
This ends the induction. The desired result is then easily obtained from the proven bound. ∎
@@
Let denote $n_{0}=|V(G_{0})|$. According to 3.3, the complexity to generate the set $\mathcal{I}$ is in $2^{\mathcal{O}\left(r\log k_{0}\right)}n_{0}^{\mathcal{O}(1)}(2r)^{\frac{k_{0}}{r-1}}$ and $|\mathcal{I}|=\mathcal{O}\left((2r)^{\frac{k_{0}}{r-1}}\right)$. Solving $(r,\mathcal{G})$-Ann-FVS on an instance $(G,k,M,\mathcal{H})\in\mathcal{I}$ with $G$ a $n$-vertex graph can be done in time $T_{DP}(k_{0},r,t)(2t)^{\frac{2k_{0}}{t-1}}n_{0}^{\mathcal{O}(1)}$, where $T_{DP}(k_{0},r,t)=2^{\mathcal{O}\left(k_{0}^{\delta}p_{7}(r,t)\right)}$ according to 4.3, 4.2 and the observation that $k\leq k_{0}$ and $n\leq n_{0}$.
@@
Now lets recall that $t$ is defined as a function of $r$ as $t=2d(r)=r^{\mathcal{O}(1)}$ and so $p_{7}(r,t)=r^{\mathcal{O}(1)}$. More precisely if we have $f_{1}(r)=\operatorname{\tilde{\mathcal{O}}}\left(r^{c_{f_{1}}}\right),~f_{2}(r,p,m)=\operatorname{\tilde{\mathcal{O}}}(r^{c_{f_{2}}}(p+m)^{c_{f_{2}}^{\prime}})$, $\tau(r)=\operatorname{\tilde{\mathcal{O}}}(r^{c_{f}})$ and $d(r)=\operatorname{\tilde{\mathcal{O}}}(r^{c_{d}})$ (remember that without loss of generality we supposed $c_{d}\geq 1$), then we have $p_{7}(r,t)=\operatorname{\tilde{\mathcal{O}}}(r^{c_{7}})$ with
@@
So by taking $r=k_{0}^{\varepsilon}$ with $\varepsilon=\frac{1-\delta}{c_{7}+1}$ we obtain a running time $2^{\operatorname{\tilde{\mathcal{O}}}\left(k_{0}^{1-\varepsilon}\right)}n_{0}^{\mathcal{O}(1)}$. To get rid of the hidden logarithmic factors in the exponent and obtain the wanted result it then suffices to replace $\varepsilon$ with some $0<\varepsilon^{\prime}<\varepsilon$. ∎